# The EarlyBird Gets the WORM: Heuristically Accelerating EarlyBird Convergence

### v1

@@ VI-B Extending WORM to Transformers
TABLE VI: Performance of WORM on GLUE Tasks
@@
As shown above, WORM achieved smaller accuracy drops than naive search but with less retraining gain and lower speedup compared to CNNs. Interestingly, WORM slowed down EB Train convergence for transformers. This might be due to distorting pre-trained information by truncating gradients. The loss plot of a run of WORM on the SST-2 task reveals this fact to be true, as the loss curve flattens very prematurely when WORM is applied.
@@
VII Conclusion
EB Train efficiently finds winning tickets but does not leverage all of the information gained during the search phase. WORM leverages this information by applying Gradient Truncation, forcing the model to learn mostly using important connections before pruning. This improves post-pruning accuracy, retraining speed, and winning ticket convergence for CNNs. While slower for transformers, WORM remains a promising first step for efficient transformer pruning methods. In the future, we hope to further investigate more complex approaches for hinting pretrained transformers, as WORM typically incites a noticeable increase in compute time, for marginally improved robustness. In addition, investigating how WORM interacts with more complex pruning strategies, such as gradient pruning, structured pruning, and undecayed pruning, can provide insight into how the behavior of prune masks generalize near the convergence boundary.
Acknowledgment
Thank you to Dr. Celine Lin and the three teaching assistants (Yonggan Fu, Haoran You, and Zhongzhi Yu) for the course CS 8803: Efficient Machine Learning at Georgia Tech, for both introducing me to the idea of Early Bird pruning, and for motivating the core idea behind this project. In addition, thank you to PACE @ Georgia Tech for providing the compute resources, to train the models involved for the project.

### v2

@@ 6.2 Extending WORM to Transformers
Table 6: Performance of WORM on GLUE Tasks
@@
As shown above, WORM achieved smaller accuracy drops than EB Train but with less retraining gain and lower speedup compared to CNNs. Interestingly, WORM slowed down EB Train convergence for transformers. This might be due to distorting pre-trained information by truncating gradients. The loss plot of a run of WORM on the SST-2 task reveals this fact to be true, as the loss curve flattens very prematurely when WORM is applied. However, more experimental analysis is needed to validate this claim.
@@
These results, taken collectively, show that directly applying WORM to transformers, while a good foundation for more intricate algorithms, does not retain ALL of the benefits it presents with convolutional neural networks. Although the experiments demonstrate that WORM enables models to retain more information even with high pruning rates (as shown by the low drops in accuracy with pruning), the additional computation overhead needed for the algorithm, combined with the matched performance in accuracy recovery may imply that this algorithm will not scale as well as it does for convolutional neural networks.
7 Conclusion
EB Train efficiently finds winning tickets but does not leverage all of the information gained during the search phase. WORM leverages this information by applying Gradient Truncation, forcing the model to learn mostly using important connections before pruning. This improves post-pruning accuracy, retraining speed, and winning ticket convergence for CNNs. While slower for transformers, WORM remains a promising first step for efficient transformer pruning methods.
In addition, despite these highly promising findings, testing WORM’s capabilities across an even more diverse set of architectures, such as RNNs and larger transformers, is quite necessary, but due to time and compute constraints, such a diverse variety of experiments was not reasonably feasible. Nevertheless, due to their performance on CNNs, WORM still remains both a robust improvement over the naive EB Train algorithm, and a great baseline for designing more complex algorithms that are more computationally efficient.
In the future, we hope to further investigate more complex approaches for hinting pretrained transformers, as WORM typically incites a noticeable increase in compute time, for marginally improved robustness. Finally, investigating how WORM interacts with more complex pruning strategies, such as gradient pruning, structured pruning, and undecayed pruning, can provide insight into how the behavior of prune masks generalize near the convergence boundary.
Acknowledgments
Thank you to Dr. Celine Lin and the three teaching assistants (Yonggan Fu, Haoran You, and Zhongzhi Yu) for the course CS 8803: Efficient Machine Learning at Georgia Tech, for both introducing me to the idea of Early Bird pruning, for motivating the core idea behind this project, and for providing feedback throughout my investigation. In addition, thank you to PACE @ Georgia Tech for providing the compute resources, to train the models needed for the project.